OpenStudy (anonymous):

Adena has a number cube whose sides are labeled 1 through 6. She rolls the number cube three times. What is the probability that Adena will roll 2, then 3, and then 4?

OpenStudy (anonymous):

1/6*1/6*1/6=1/216

Mehek (mehek14):

the total is 6 rolling a 2 \(\dfrac{1}{6}\) rolling 3 \(\dfrac{1}{6}\) rolling 4 \(\dfrac{1}{6}\) all you have to do is multiply \(\dfrac{1}{6}*\dfrac{1}{6}*\dfrac{1}{6}\)

Mehek (mehek14):

which gives you \(\dfrac{1}{216}\)

OpenStudy (anonymous):

Simplified?

OpenStudy (anonymous):

\[\left(\frac{1}{6}\right)^3=\frac{1}{216} \]

OpenStudy (loser66):

@robtobey I have a question on it: For the first time we roll the cube, the probability of number 2 is 1/6 but, how can we calculate the probability of number 2 which happen right at the first time we roll?? Likely, for number3 for the second time and number 4 for the last one? I know they are independent.

OpenStudy (loser66):

I meant: if we roll the cube 6 times, 2 can happen once. How to know that 2 will happen the first time we roll?

OpenStudy (loser66):

If I interpret the problem like this, is it valued? We have the probability of number 3 is 1/6, hence the first time rolling, we need the sample is NOT3, hence the probability of NOT3 is 5/6 same as NOT4, NOT5, NOT6 and NOT1, add them together, If the first time we roll is NOT 1,3,4,5,6 , that force the first time is 2 and the probability to get the first time is number 2 is .....

OpenStudy (loser66):

* not "add" , multiply them.

OpenStudy (anonymous):

The outcomes are sequential. First a 2 must appear on the first roll, then a 3 on the second roll and finally a 4 on the third roll. The odds of rolling a 2 and then a 3 in that order is (1/6)*(1/6) . The odds of rolling a 2, 3 and then 4, the required ordered outcome, is (1/6)*(1/6)*(1/6).

OpenStudy (anonymous):

By the way, the odds of failing to roll 2, 3 and 4 in that order is:\[1-\left(\frac{1}{6}\right)^3=\frac{215}{216} \],

OpenStudy (loser66):

I need time to digest it. Thanks for explanation.